### v1

# Self-regularized entropy: How much do quantum black holes violate the no-hair theorem?

Abstract
We compute the canonical (brick wall) entropy of Hawking radiation in a quantum black hole whose exterior is described, to first order in a small quadrupole parameter, by the static $q$-metric, which is an exact vacuum solution of the Einstein equations. Counting near horizon quasinormal modes shows that a modest quadrupolar deformation self-regularizes the ultraviolet divergence: the entropy of Hawking radiation is finite for any non-vanishing quadrupole, without an ad hoc cutoff. Matching this canonical entropy to the Bekenstein-Hawking entropy leads to no-hair violating multipoles, at percent-to-tens-of-percent level, and provides concrete observational targets for the Next Generation Event Horizon Telescope (ngEHT) and the Laser Interferometer Space Antenna (LISA).
preprint: APS/123-QED
I Introduction
Despite the success of black hole thermodynamics, the microscopic origin of the Bekenstein-Hawking entropy remains unsettled; a variety of horizon degrees of freedom have been proposed e.g., PhysRevD.7.2333 ; 1998PhRvL..81.4293H ; 1996PhLB..379…99S ; 1997NuPhS..57…65I ; 2000gr.qc…..5126A ; 2003PhRvL..90h1301D ; 2004CQGra..21.5245M ; 2006CQGra..23.3849D ; Saravani:2012is ; 2021PhRvD.104l1502C ; 2023PhLB..84137901O ; 2023arXiv230404491N , yet no single picture is universally accepted. In this work, we explore a conservative, geometry-based option: we treat the quasinormal (QN) modes as an effective set of near horizon excitations that can account for (at least part of) the entropy. The idea is to model quantum fluctuations of the spacetime geometry as superpositions of QN mode eigenstates, in close analogy with the quantization of a harmonic oscillator, where the occupation of each mode provides a natural bookkeeping of microstates.
The brick wall model of ’t Hooft tHooft:1984kcu ; tHooft:1996rdg remains the simplest way to estimate the canonical (thermodynamic) entropy of quantum fields around a black hole, but in the exactly spherical Schwarzschild geometry the mode sum diverges logarithmically and one is forced to introduce an ad‑hoc proper distance cutoff. In this paper, we show that the brick wall ultraviolet divergence disappears once the Schwarzschild exterior is perturbed by an arbitrarily small quadrupolar deformation.
An implication of a brick wall model for a quantum black hole is that we can send and receive signals from the quantum structure near the horizon, within a finite time. This has motivated searches for “echoes” following black hole mergers in gravitational wave e.g, Abedi:2016hgu ; Cardoso:2016rao ; Oshita:2019sat ; Wang:2019rcf ; Oshita:2023tlm ; Vellucci:2022hpl ; Arrechea:2024nlp (see Abedi:2020ujo ; Berti:2025hly for reviews). However, this would also imply that a stationary spacetime may not obey the no-hair theorem, as it can learn about the absence of a regular horizon in a finite time (e.g., Bah:2021jno ). In this paper, we propose that the $q$-metric may be used as a way to probe the impact of these near horizon quantum effects on the rest of the exterior spacetime, through solving the Einstein field equations.
Our goal is to isolate the single geometric ingredient that makes the Schwarzschild brick wall sum diverge and to test how a minimal, physically motivated deformation changes the near horizon density of states. The static Zipoy-Voorhees $q$-metric osti_4201189 ; PhysRevD.39.2904 is ideal for this. It is an exact, asymptotically flat, one parameter deformation of Schwarzschild that remains a vacuum solution of Einstein’s equations and reduces smoothly to Schwarzschild as $q\,\to\,0$. The single parameter $q$ encodes a quadrupolar departure from spherical symmetry; the first nontrivial hair allowed once perfect symmetry is relaxed. Working in vacuum general relativity avoids modeling ambiguities from matter sources or modified field equations and keeps all changes traceable to geometry alone.
We work with the $q$-metric line element e.g., osti_4201189 ; PhysRevD.39.2904 , a one parameter ($q$) vacuum deformation of Schwarzschild, and restrict to first order in $|q|\ll 1$. The transverse circumference pinches (for $q>0$) and the density of states softens. A direct count of quasinormal modes shows that the canonical entropy converges for a range of $q>0$; no brick wall cutoff or stretched horizon dissipation is required. Thus, the $q$-metric delivers a closed form, analytically tractable test of self-regularisation without introducing ad hoc cutoffs or dissipation. Treating $q$ as the coarse grained imprint of horizon microstates and imposing that the entropy radiated by Hawking quanta never exceeds the Bekenstein-Hawking entropy fixes its natural scale to $|q|\simeq 0.1$ for all astrophysical black holes. A distortion of this magnitude is not merely theoretical: it shifts the lowest mass and current multipole moments linearly in $q$, supplying a concrete observational target, e.g., for next‑generation horizon imaging (ngEHT) and gravitational wave spectroscopy (LISA). We also show that in the stationary extension of the $q$-metric the same parameter $q$ induces linear, Kerr-violating shifts in the lowest multipoles captured by two Kerr null combinations $Q_{1}$ and $Q_{2}$, so that the entropy motivated bound on $|q|$ becomes an observational target for ngEHT and LISA.
We emphasise that we are not claiming astrophysical objects are exact $q$-metrics. Rather, the $q$-metric is the simplest vacuum laboratory that (a) preserves the universal Rindler $\times$ transverse near horizon structure, (b) introduces precisely one controlled deformation that governs the ultraviolet behaviour of the mode sum, and (c) connects directly to measurable multipoles. More general ”bumpy” metrics or rotating spacetimes introduce additional scales (e.g. superradiance, spin couplings) that obscure this single parameter test without changing the qualitative conclusion: any axisymmetric deformation, self-regularises the brick wall in the same way; the $q$-metric is simply the cleanest analytic example.
The rest of the manuscript is orgnaized as follows: Section II reviews the $q$-metric. In Section III we derive the scalar wave operator and the exact radial Klein-Gordon problem, while Section IV performs the WKB counting and presents the entropy. Section V offers observational prospects for measuring quadrupole, and Section VI gathers the main implications. In this work, we adopt the convention $G=c=\hbar=k_{B}=1$.
II Spacetime
The $q$-metric describes static, axially symmetric, and asymptotically flat exact solutions to the vacuum Einstein field equations. By incorporating the quadrupole moment, the $q$-metric accounts for the deformation of the gravitational field which is not captured by the spherically symmetric Schwarzschild solution. Consequently, the $q$-metric characterizes the exterior gravitational field of an isolated static axisymmetric object. The presence of a quadrupole can change the geometric properties of spacetime and affect the timelike and lightlike trajectories e.g., 2021A&A…654A.100F ; 2025PhRvD.111d5003A ; 2025A&A…699A.266F ; 2025EPJC…85..148F . As a result, the $q$-metric is a particularly useful toy model for the gravitational fields of realistic astrophysical bodies where the effects of higher-order multipole moments cannot be ignored. The metric in Schwarzschild-like coordinates $(t,x,y,\phi)$ is presented as follows osti_4201189 ; PhysRevD.39.2904
$ds^{2}=$ $\left(1-\frac{2m}{r}\right)^{1+{q}}dt^{2}-\left(1-\frac{2m}{r}\right)^{-{q}}$
$\left[\left(1+\frac{m^{2}\sin^{2}\theta}{r^{2}-2mr}\right)^{-{q}(2+{q})}\left(\frac{dr^{2}}{1-\frac{2m}{r}}+r^{2}d\theta^{2}\right)\right$.
$\left.+r^{2}\sin^{2}\theta d\phi^{2}\right],\$(1)
where $t\in(-\infty,+\infty)$, $r\in(2m,+\infty)$, $\theta\in[0,\pi]$, and $\phi\in[0,2\pi)$, also $m$ is a parameter which is expressed in the dimension of length. This metric contains two free parameters, namely $m$, and quadrupole parameter $q$. In the case of vanishing $q$, the Schwarzschild metric is recovered. Furthermore, values of $q>0$ describe the exterior field of an oblate central object, while $q<0$ is related to a prolate source. Additionally, the $q$-metric, aside from a central curvature singularity at $r=0$, has another singularity located at a finite distance from the origin at $r=2m$ for any chosen nonzero value of quadrupole. By Geroch’s definition (1970JMP….11.2580G, ), for multipole moment and for avoiding a negative mass distribution, the quadrupole parameter is restricted to the domain of $q\in(-1,\infty)$.
II.1 Echo time and entropy bound
The time scale for a null signal to reach the Schwarzschild radius $r=2m$ from the light ring $r=3m$ and back, in the static $q$-metric is given by
$t_{\rm echo}$ $\simeq 2\int_{2m}^{3m}dr\left(1-\frac{2m}{r}\right)^{-1-q}\left(\frac{r^{2}-2mr+m^{2}\sin^{2}\theta}{r^{2}-2mr}\right)^{-\frac{q(2+q)}{2}}$ (2)
$\simeq 8mq^{-2}+{\cal O}(mq^{-1})=\frac{1}{\pi T_{H}q^{2}}.$
We take the asymptotic Hawking temperature to be that of a Schwarzschild black hole with the same mass, $T_{H}=1/(8\pi m)$. This is a modeling choice consistent with the small-$q$ limit and with our use of the near horizon cavity to count modes Oshita_2023 . The $\theta$ dependence in the integrand produces only an $\mathcal{O}(m)$ modulation and does not affect the leading $mq^{-2}$ scaling  [^1] To first order in $q$, the light-ring radius shifts to $r_{\rm LR}=m(3+2q)$; using $3m$ as the upper limit changes $t_{\rm echo}$ only at $\mathcal{O}(q^{0})$, hence subleading to the $q^{-2}$ term.. We note that for the entropy in the Hawking radiation not to exceed the Bekenstein-Hawking horizon entropy of the black hole Oshita_2023 , we require
$t_{\rm echo}\lesssim\frac{\ln S}{2\pi T_{H}}\Rightarrow q^{2}\gtrsim\frac{2}{\ln S}.$ (3)
This implies that, while any value of $q$ solves the classical vacuum Einstein equations, the finite entropy of quantum effects that lead to Hawking radiation puts a lower bound on $|q|$, that would describe a semi-classical spacetime. This sets the minimal modifications we expect from Schwarzschild/Kerr solutions, even far from the horizon.
Equation (3) should be read as a conservative lower bound on $q$. It is obtained by treating the Hawking flux and surface gravity as those of Schwarzschild (i.e. $T_{H}=1/8\pi M$) and by using the leading small $|q|$ echo time scaling, so that the total entropy trapped within the angular momentum barrier does not exceed the Bekenstein-Hawking value Oshita_2023 . This estimate merely sets the minimal distortion required to avoid overcounting. In what follows (Sections III-V), we compute the canonical entropy directly in the $q$-metric by near horizon WKB mode counting, which yields a sharper relation $q=q(S)$ for this spacetime and thus a more accurate bound or estimate on $q$.
II.2 Small-$q\ll 1$ near horizon limit
In this part, we derive the near horizon line element and the associated tortoise coordinate for the static $q$-metric (II). Throughout, we keep the leading non-vanishing power of the affine distance to the horizon and work in $|q|\ll 1$ limit. This computation will prepare us for the WKB mode counting in the next section, which shall yield a finite enetropy of Hawking radiation in the q-metric spacetime.
In the limit $q\ll 1$, the metric simplifies and the exponents in $\rho$ become approximately symmetric. Introducing the dimensionless parameter $\hat{\rho}=1-2m/r$, so that $\hat{\rho}\to 0^{+}$ as $r\to 2m$. We then pass to the Gaussian-normal coordinate $\rho$ by
$\rho=\hat{C}\,\hat{\rho}^{\,\beta},\quad\text{where}\quad\beta:=\frac{q^{2}}{1+q+q^{2}},$ (4)
with an arbitrary constant scale $\hat{C}$ (irrelevant hereafter). In these coordinates, this 2D $(t,\rho)$ sector takes a Rindler-like form $f(\rho)dt^{2}+d\rho^{2}$, with
$f(\rho)=\left(\frac{\rho}{4m}\right)^{\alpha},\qquad\alpha:=2(1-\beta)=\frac{2(1+q)}{1+q+q^{2}},$ (5)
approaching Rindler in $q\to 0\Rightarrow\alpha\to 2$ limit. The angular 2-metric is a mild deformation of the round sphere. To leading order in $|q|$ the full near horizon line element is
$ds^{2}$ $=f(\rho)\left(-dt^{2}+\frac{1}{f(\rho)}d\rho^{2}\right)$
$+4m^{2}\left[\left(\frac{\rho}{4m}\right)^{2q}d\theta^{2}+\left(\frac{\rho}{4m}\right)^{-2q}\sin^{2}\theta\,d\phi^{2}\right].$ (6)
This procedure is equivalent to zooming into a local Rindler patch, but it preserves the qq‑dependent exponents that control the ultraviolet behaviour of the mode sum .
We can consider spacetime in equation (II.2) as a generalized Rindler patch, times a transverse 2-manifold. We are using this Rindler form of metric to simplify the computation of entropy in the WKB approximation. In the high frequency limit, a field mode is a localized wave packet whose phase advances according to the null Hamilton-Jacobi relation $g^{\mu\nu}p_{\mu}p_{\nu}=0$. The number of modes with frequency below $\omega$ is, to leading order, the phase‑space volume on a constant time slice divided by $(2\pi)^{3}$: $N(\omega)\ \sim\ \frac{1}{(2\pi)^{3}}\,\int d^{3}x\,d^{3}p\,\Theta\,\bigl{[}\omega-\mathcal{H}\,(x,p)\bigr{]}$ where $\mathcal{H}$ is determined by the local metric. Therefore, only two local inputs control the ultraviolet behavior: (a) the redshift $g_{tt}=f(\rho)$ in the radial Rindler sector, and (b) the intrinsic metric of the transverse 2-surface (which fixes the map between angular quantum numbers and transverse momenta). For near horizon Schwarzschild, $g_{tt}\propto\rho^{2}$ and the transverse radius $r(\rho)$ tends to a constant, $2m$. The local phase space measure near the horizon contains $d\rho/\rho$ - a logarithmically divergent integral- since the redshift creates arbitrarily many radial “boxes” of size set by the local wavelength. In the static $q$-metric, the transverse circumference is not constant as in equation (II.2). This single geometric change modifies the phase space measure by a factor $\rho^{\,\pm 2q}$ coming from the transverse area and the centrifugal barrier. These two ingredients are precisely what the deformed Rindler $\times$ (deformed $S^{2}$) neighborhood encodes, and they are exactly what turn the divergent Schwarzschild log into a convergent power when $|q|>0$ as we see in the rest of the paper.
II.2.1 Tortoise Coordinates
We define the tortoise coordinate $r^{*}$ in the usual way, so that the radial part of the Klein-Gordon operator has the Schrödinger form, and integrating gives
$dr_{*}$ $=\frac{d\rho}{\sqrt{f(\rho)}}=(4m)^{\alpha/2}\,\rho^{-\alpha/2}\,d\rho$
$\Rightarrow\quad r_{*}=\frac{(4m)^{\alpha/2}}{\beta}\,\rho^{\,\beta},$ (7)
where $\beta$ is given by equation (4). Inverting $\rho=\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!1/\beta}$, turns the metric (II.2) to
$ds^{2}$ $=-(4m)^{-\alpha}\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!\alpha/\beta}\,dt^{2}+(4m)^{-\alpha}\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!\alpha/\beta}\,dr_{*}^{2}$
$\quad+4m^{2}\left[(4m)^{-2q}\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!2q/\beta}d\theta^{2}\right$.
$\quad\left.+(4m)^{\,2q}\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!-2q/\beta}\sin^{2}\theta\,d\phi^{2}\right]\,.$ (8)
III Wave Operator Setup
For the massless scalar field the equation of motion is
$\Box\,\varphi\;=\;\frac{1}{\sqrt{-g}}\,\partial_{\mu}\,\bigl{(}\sqrt{-g}\,g^{\mu\nu}\partial_{\nu}\varphi\bigr{)}=0.$ (9)
Using the near horizon metric (II.2) one obtains $\sqrt{-g}=4m^{2}\sqrt{f}\sin\theta$. A component by component evaluation (see appendix A) of equation (9) gives
$\Box\varphi$ $\,=-\frac{1}{f(\rho)}\,\partial_{t}^{2}\varphi+\partial_{\rho}^{2}\varphi+\frac{1}{2}\,\frac{f^{\prime}(\rho)}{f(\rho)}\,\partial_{\rho}\varphi$ (10)
$+\frac{1}{4m^{2}}\left[\left(\frac{\rho}{4m}\right)^{-2q}\frac{1}{\sin\theta}\,\partial_{\theta}\!\left(\sin\theta\,\partial_{\theta}\varphi\right)+\left(\frac{\rho}{4m}\right)^{2q}\frac{1}{\sin^{2}\theta}\,\partial_{\phi}^{2}\varphi\right].$
III.1 Separation of variables
To obtain a single radial equation we use the standard mode ansatz
$\varphi(t,\rho,\theta,\phi)=e^{-i\omega t}\,Y_{\ell m}(\theta,\phi)\,\psi(\rho),\quad\Delta_{S^{2}}Y_{\ell m}=-\ell(\ell+1)Y_{\ell m}.$ (11)
Because the angular operator carries unequal $\rho^{\mp 2q}$ weights in the $\theta$ and $\phi$ parts, exact separation into a single ODE holds rigorously in the axisymmetric sector $m=0$. However, for general $m$ and $|q|\ll 1$ one may use the $\ell$-dominated approximation. Replacing the weighted angular operator by its $\ell(\ell+1)$ eigenvalue; this is accurate at leading order in $q$ (the corrections are $O(q\,/\ell)$ and do not affect the near horizon exponents or the WKB scaling).
Projecting (10) onto $Y_{\ell m}$ yields the radial equation
$\psi^{\prime\prime}(\rho)+\frac{1}{2}\frac{f^{\prime}(\rho)}{f(\rho)}\,\psi^{\prime}(\rho)+\left[\frac{\omega^{2}}{f(\rho)}-\frac{\ell(\ell+1)}{r^{2}(\rho)}\right]\psi(\rho)=0,$ (12)
where
$r^{2}(\rho)=4m^{2}\left(\frac{\rho}{4m}\right)^{2q}.$ (13)
To remove the first derivative term we pass to tortoise coordinates (7), $dr_{*}=d\rho/\sqrt{f(\rho)}$, and perform the Liouville transform $\Psi(\rho):=f(\rho)^{-1/4}\psi(\rho)$. Writing $\Psi(r_{*})=\Psi(\rho(r_{*}))$ we obtain the Schrödinger form
$\frac{d^{2}\Psi}{dr_{*}^{2}}+\bigl{[}\omega^{2}-V_{\ell}(\rho)\bigr{]}\Psi=0,$ (14)
where
$V_{\ell}(\rho)=f(\rho)\,\left[\frac{\ell(\ell+1)}{r^{2}(\rho)}+\frac{1}{4}\,\left(\frac{f^{\prime}(\rho)}{f(\rho)}\right)^{2}-\frac{1}{2}\,\frac{f^{\prime\prime}(\rho)}{f(\rho)}\right].$ (15)
For our $f(\rho)=\bigl{(}\rho/(4m)\bigr{)}^{\alpha}$ one has $f^{\prime}/f=\alpha/\rho$ and $f^{\prime\prime}/f=\alpha(\alpha-1)/\rho^{2}$, hence simplifies the bracket
$\frac{1}{4}\Bigl{(}\frac{f^{\prime}}{f}\Bigr{)}^{2}-\frac{1}{2}\frac{f^{\prime\prime}}{f}=\frac{\alpha\beta}{2\,\rho^{2}},$ (16)
Therefore, we obtain
$V_{\ell}(\rho)=\frac{\ell(\ell+1)}{4m^{2}}\left(\frac{\rho}{4m}\right)^{\alpha-2q}\,+\;\frac{\alpha\beta}{32\,m^{2}}\left(\frac{\rho}{4m}\right)^{\alpha-2}\,.$ (17)
In tortoise coordinates.
Using $\rho=\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!1/\beta}$ and noting $\alpha-2=-2\beta$, the curvature piece becomes a universal inverse square wall,
$\frac{\alpha\beta}{32\,m^{2}}\left(\frac{\rho}{4m}\right)^{\alpha-2}=\frac{\alpha}{2\beta}\,\frac{1}{r_{*}^{2}},$ (18)
and the centrifugal term scales as
$\frac{\ell(\ell+1)}{4m^{2}}$ $\left(\frac{\rho}{4m}\right)^{\alpha-2q}=\frac{\ell(\ell+1)}{4m^{2}}\left(\frac{\beta\,r_{*}}{4m}\right)^{\!p},$ (19)
where
$p:=\frac{\alpha-2q}{\beta}=\frac{2}{q^{2}}-2-2q.$ (20)
Thus, the master equation becomes
$\frac{d^{2}\Psi}{dr_{*}^{2}}+\Bigl{[}\omega^{2}-V_{\ell}(r_{*})\Bigr{]}\Psi=0,$ (21)
where
$V_{\ell}(r_{*})=\frac{\ell(\ell+1)}{4m^{2}}\left(\frac{\beta\,r_{*}}{4m}\right)^{\!p}+\frac{\alpha}{2\beta}\,\frac{1}{r_{*}^{2}}.$ (22)
Equation (21) is the precise radial master equation that enters the WKB mode counting. All powers and coefficients follow directly from the metric exponents $\alpha$ and $\beta$.
III.2 Classical turning points and cavity length
The WKB approximation is valid in the domain where the radial wave number $k_{r}^{2}\,=\,\omega^{2}-V_{\ell}(\rho)$ is positive, i.e. between two classical turning points. For the potential $V_{\ell}(\rho)$ in equation (17) the inner edge is $\rho_{\min}=0$, while the outer edge $\rho_{\max}$ is defined by $k_{r}^{2}(\rho_{\max})=0$:
Outer turning point.
Using $f(\rho)=\bigl{(}\rho/(4m)\bigr{)}^{\alpha}$ and $r^{2}(\rho)=4m^{2}\bigl{(}\rho/(4m)\bigr{)}^{2q}$, the transverse wavenumber is
$k_{\perp}^{2}(\rho):=\frac{\ell(\ell+1)}{r^{2}(\rho)}\simeq\frac{\ell^{2}}{4m^{2}}\left(\frac{\rho}{4m}\right)^{-2q},$ (23)
(valid for  $\ell\,\gg\,1$). Setting $\omega^{2}=f(\rho_{\max})\,k_{\perp}^{2}(\rho_{\max})$ gives
$\left(\frac{\rho_{\max}}{4m}\right)^{\alpha-2q}$ $=\frac{4m^{2}\,\omega^{2}}{\ell^{2}}$
$\Longrightarrow\quad\rho_{\max}=4m\left(\frac{4m^{2}\omega^{2}}{\ell^{2}}\right)^{\!1/(\alpha-2q)}.$ (24)
Tortoise position of the turning point.
With equation (7), $r_{*}=\dfrac{(4m)^{\alpha/2}}{\beta}\,\rho^{\beta}=\dfrac{4m}{\beta}\left(\dfrac{\rho}{4m}\right)^{\beta}$, we obtain
$r^{*}_{\max}=\frac{4m}{\beta}\,\left(\frac{4m^{2}\omega^{2}}{\ell^{2}}\right)^{\!\beta/(\alpha-2q)},\left(\approx\frac{4m}{q^{2}}\left(\frac{2m\omega}{\ell}\right)^{q^{2}}\right)$ (25)
which approaches $t_{\rm echo}/2$ (equation 2), in the $q\ll 1$ limit.
Cavity length.
Since the inner edge is at $r^{*}_{\min}=0$, the proper WKB cavity length is simply
$L(\omega,\ell)=r^{*}_{\max}.$ (26)
Inner turning point (why one may set $\rho_{\min}=0$).
Besides the centrifugal term, the effective potential contains the “curvature” part
$V_{\mathrm{curv}}(\rho)=f(\rho)\,\frac{\alpha\beta}{2\rho^{2}}=\frac{\alpha\beta}{32\,m^{2}}\left(\frac{\rho}{4m}\right)^{\alpha-2},\quad 0<\alpha<2,\ \beta>0,$ (27)
where we used $f(\rho)=\bigl{(}\rho/(4m)\bigr{)}^{\alpha}$ and $\alpha-2=-2\beta$. Since $\alpha<2$ for every $|q|<1$, one has $V_{\mathrm{curv}}\to+\infty$ as $\rho\to 0$, so the inner (classical) turning point is determined by
$\omega^{2}=V_{\mathrm{curv}}(\rho_{\mathrm{in}})$ $\;\;\Longrightarrow\;\;\left(\frac{\rho_{\mathrm{in}}}{4m}\right)^{\alpha-2}=\frac{32\,m^{2}\,\omega^{2}}{\alpha\beta}$
$\;\;\Longrightarrow\;\;\rho_{\mathrm{in}}=4m\left(\frac{\alpha\beta}{32\,m^{2}\,\omega^{2}}\right)^{\!1/(2-\alpha)}.$ (28)
In tortoise coordinates $r^{*}=\dfrac{4m}{\beta}\left(\dfrac{\rho}{4m}\right)^{\beta}$, this maps to
$r^{*}_{\mathrm{min}}=\frac{4m}{\beta}\left(\frac{\rho_{\mathrm{in}}}{4m}\right)^{\beta}=\frac{4m}{\beta}\left(\frac{\alpha\beta}{32\,m^{2}\,\omega^{2}}\right)^{\!1/2}=\frac{1}{\omega}\,\sqrt{\frac{\alpha}{2\beta}}\approx\frac{1}{\omega q}\,$,
which tends to zero for large frequencies. Furthermore, $r^{*}_{\mathrm{min}}\ll r^{*}_{\mathrm{max}}$ for $q\ll 1$ and $\omega\sim T_{\rm H}$. Thus the divergent curvature term acts effectively as a hard wall at $r^{*}\approx 0$. Replacing $\rho_{\mathrm{in}}$ by 0 therefore shifts the cavity length $L$ only by an $\mathcal{O}(q)$ fraction and does not affect the WKB scaling or the finiteness of the entropy. This justifies setting $\rho_{\min}=0$ in equation (26).
For $q=0$ (Schwarzschild) equation (25) reduces to the familiar logarithmic behavior $L\sim 4m\ln\big(\ell/(m\omega)\big{missing})$+ const. For any $|q|\ll 1$, it becomes a power law in $\omega/\ell$ displayed by equation (25); this change of scaling underlies the self-regularisation of the mode sum discussed in the next section.
IV Mode counting and canonical entropy
In this section we derive the entropy of a single massless scalar field within the $q$-metric, at Hawking temperature.
IV.1 WKB radial spectrum
For each $(\ell,m)$ the radial wave‑number is $k_{r}(\rho)=\sqrt{\omega^{2}-V_{\ell}(\rho)}$. With $r_{*}^{\min}=0$ and $r_{*}^{\max}=L$ (Sec. III.2). In the WKB approximation, for fixed $(\ell,m)$ the counting function $N_{\ell}(\omega)$ is the radial mode number with frequency $\leq\omega:=\omega_{n\ell}$ is smooth and given by
$n:=N_{\ell}(\omega)=\frac{1}{\pi}\int_{0}^{L}\,k_{r}\,dr_{*},$ (29)
inside the cavity we have $k_{r}\simeq\omega$, therefore $n\simeq\frac{\omega L(\omega,\ell)}{\pi}$ and from equations (25) and (26) we have
$L(\omega,\ell)=\frac{4m}{\beta}\,\left(\frac{4m^{2}\omega^{2}}{\ell^{2}}\right)^{\!\beta/(\alpha-2q)}.$ (30)
Writing $L(\omega,\ell)=C(q,m)\,\ell^{-s}\omega^{s}$ gives the level law $\omega:=\omega_{n\ell}\ \propto\ n^{\delta}\,\ell^{\sigma}$ with $\delta:=\frac{1}{1+s},\sigma:=\frac{s}{1+s}$, with
$\quad s:=\frac{2}{p}=\frac{2\beta}{\alpha-2q}=\frac{q^{2}}{1-q^{2}-q^{3}}>0,$ (31)
and
$C(q,m)=\frac{4m}{\beta}\,(4m^{2})^{\,\beta/(\alpha-2q)}=\frac{4m}{\beta}\,(4m^{2})^{\,s/2}.$ (32)
IV.2 Radial density of states
By defining the dimensionless frequency $w=\omega/T_{H}$, the canonical entropy of one bosonic oscillator is $\mathcal{S}(w)=\dfrac{w}{e^{w}-1}-\ln\,\bigl{(}1-e^{-w}\bigr{)}$. The total entropy is then given by a discrete sum over all the modes:
$S(q)$ $\equiv\sum^{\infty}_{\ell=0}(2\ell+1)\sum_{n=1}^{\infty}\mathcal{S}\left(\frac{\omega_{n\ell}}{T_{H}}\right)\simeq\int_{0}^{\infty}2\ell d\ell\sum_{n=1}^{\infty}\mathcal{S}(w)$
$=\int_{0}^{\infty}\left(\sum_{n=1}^{\infty}2\ell\frac{\partial\ell(w,n)}{\partial w}\right)\mathcal{S}(w)dw,$ (33)
with $n=\left(\frac{C}{\pi}\right)\ell^{-s}\omega^{s+1}$, and differentiate at fixed $n$ finds
$\sum_{n=1}^{\infty}2\ell\,\frac{\partial\ell}{\partial w}=\frac{2(1+s)}{s}\,\left(\frac{C}{\pi}\right)^{\frac{2}{s}}\,\zeta\,\left(\frac{2}{s}\right)\,T_{H}^{\frac{2(1+s)}{s}}w^{\frac{2(1+s)}{s}-1}.$ (34)
Therefore
$S(q)=\frac{2(1+s)}{s}\,\left(\frac{C}{\pi}\right)^{\frac{2}{s}}\,\zeta\,\left(\frac{2}{s}\right)\,T_{H}^{\frac{2(1+s)}{s}}\int_{0}^{\infty}w^{\frac{2}{s}+1}\,\mathcal{S}(w)\,dw,$
and using
$\int_{0}^{\infty}w^{u}\,\mathcal{S}(w)\,\mathrm{d}w=(u+2)\,\Gamma(u+1)\,\zeta(u+2),\quad(u>-1)$
with $u=p+1$ and $p=\frac{2}{s}$ we obtain
$S(q)=(p+2)\,\left(\frac{C}{\pi}\right)^{p}\,T_{H}^{p+2}\,\zeta(p)(p+3)\,\Gamma(p+2)\,\zeta(p+3).$
From the turning point analysis equation (26) and equation (30), and recall that $L(\omega,\ell)=C(q,m)\,\ell^{-s}\omega^{s}$ we can write
$\frac{C}{\pi}\,=\frac{4m}{\beta\pi}(4m^{2})^{s/2},\,$ (35)
using $T_{H}=(8\pi m)^{-1}$ and $p=\frac{2}{s}$ we find
$\left(\frac{C}{\pi}\right)^{p}\,T_{H}^{p+2}=(4m^{2})\left(\frac{4m}{\beta\pi}\right)^{p}(8\pi m)^{-(p+2)}$ $=2^{-(p+4)}\,\beta^{-p}\,\pi^{-(2p+2)},\,$
the factors of $m$ canceled out. Thus, the final result is
$S(q)=\frac{(p+2)(p+3)}{2^{\,p+4}\,\pi^{\,2p+2}}\;\beta^{-p}\;\Gamma(p+2)\,\zeta(p)\,\zeta(p+3),$ (36)
where $p\approx\frac{2}{q^{2}}$ and $\beta\approx q^{2}$ from equations (20) and (4). This entropy is finite for all sufficiently small $|q|\neq 0$. In contrast, as $q\to 0$ one has $p\to\infty$ (hence $s(q)\to 0^{+}$), the cavity length reverts to the Schwarzschild logarithm, and the entropy reproduces the usual brick wall divergence. In the formal zeta-regularized notation used earlier, this appears as the symbolic limit $\bigl{[}2^{-3}/q^{4}\bigr{]}\Gamma(0)\,\zeta(1)$ when $q\to 0^{+}$.
V Observational prospects
For any four‑dimensional, non‑extremal black hole
$S_{\rm BH}=\frac{A_{H}}{4G\hbar}\sim 4\pi\,\Bigl{(}\frac{m}{m_{\mathrm{P}}}\Bigr{)}^{2},$ (37)
where $m_{\mathrm{P}}=2.18\times 10^{-5}\,\mathrm{g}$, is the Planck mass  [^2] For nonextremal Kerr, $8\pi M^{2}\leq A_{H}\leq 16\pi M^{2}$ in $G=c=\hbar=1$, so $\ln S_{\rm BH}$ changes by at most $\ln 2$, negligible for our purposes.. Therefore, $\ln S_{\rm BH}$ grows only logarithmically with the mass. Using the bound (3) we set earlier in Section V.1,
$|q|\;\gtrsim\;\sqrt{\frac{2}{\ln S_{\rm BH}}}\;,$ (38)
we find that across the entire stellar to supermassive interval $5\,M_{\odot}\leq M\leq 10^{10}\,M_{\odot}$, leading to $S_{\rm BH}\,\sim\,10^{78}\text{--}10^{97}$, the bound varies only from $0.095$ to $0.11$ (light stellar) to $0.095$ (most massive).
However, from the entropy that we derived in the previous section, equation (36), we can put a stronger bound on $q$. For the astrophysically relevant small $|q|$ (and hence large $p$), by applying $\zeta(p+k)\to 1$, and Stirling’s formula for $\Gamma(p+2)$, thus $S(q)$ reduces to $p\log p$ relation which we invert with the product-log (Lambert) function $W$, defined by $W(z)\,e^{W(z)}=z$ for $z>0$, and we obtain
$q(S)\approx\sqrt{\frac{4W\left(\frac{\ln S}{4\pi\sqrt{e}}\right)}{\ln S}}.$ (39)
Plugging $S_{BH}\simeq 10^{78}-10^{97}$, i.e., $\ln S\simeq 180-223$ into the estimated formula above we obtain the more accurate bound on $q$ as
$|q|_{\rm new}\;\simeq\;0.18\,-\,0.19,$ (40)
which varies by less than 5% across 10 orders of magnitude in mass, for the black hole entropies across the stellar to supermassive mass range. This weak dependence on entropy is a strength: it means that $10-20\%$ quantum corrections to Kerr geometry are essentially mass independent and set by the absolute value of $|q|\sim 0.2$, while the spacetime remains close to Schwarzschild beyond a few gravitational radii.
V.1 From static to stationary: Multipoles
The analysis in this paper has deliberately focused on the static $q$-metric, which is the only input needed to establish the near horizon self-regularisation of the canonical entropy. However, in this section we briefly connect that static description to stationary (spinning) black hole observables, solely to indicate how a nonzero $q$ would manifest as controlled deviations from the Kerr multipole hierarchy; here we use it only as a phenomenological map from $q$ to measurable multipoles. The $q$-metric admits a stationary (spinning) extension in which the mass and current multipole moments $(M_{\ell},J_{\ell})$ depart perturbatively from the Kerr values $M_{\ell}+iJ_{\ell}=M(ia)^{\ell}$. To lowest orders one finds (e.g., Toktarbay_2014 ; Faraji_2022 ), namely we have
$M_{0}=m+q\sigma,$
$M_{2}=-m^{3}-3m^{2}q\sigma-m\left(q^{2}-1\right)\sigma^{2}-\frac{1}{3}q\left(q^{2}-7\right)\sigma^{3},$
$J_{1}=a(m+2q\sigma),$
$J_{3}=-\frac{a}{3}\left[3m^{3}+12m^{2}q\sigma+3m\left(3q^{2}-1\right)\sigma^{2}+2q\left(q^{2}-4\right)\sigma^{3}\right],$
for the lowest mass and angular momentum multipoles, where $\sigma=\sqrt{m^{2}-a^{2}}$ (e.g., Toktarbay_2014 ; Faraji_2022 ).
We define two dimensionless combinations of these mutipoles that vanish identically for Kerr
$Q_{1}\equiv\frac{|J_{1}^{2}+M_{0}M_{2}|}{M_{0}^{4}}=\frac{2}{3}\left(1-a_{*}^{2}\right)^{3/2}|q|+{\cal O}(q^{2}),$ (41)
$Q_{2}\equiv\frac{|J_{1}^{3}+M_{0}^{2}J_{3}|}{M^{5}_{0}}=\frac{4}{3}|a_{*}|\left(1-a_{*}^{2}\right)^{3/2}|q|+{\cal O}(q^{2}),$ (42)
where $a_{*}\equiv a/m$ is the dimensionless spin parameter, which ranges between 1-1 and $1$ for Kerr geometry (with $a_{*}=0$ for Schwarzschild). We see that violations of no-hair theorem are encoded in $(Q_{1},Q_{2})$, thus probing the potential quantum corrections induced by a finite value $q$.
But what value of $q$ should we use for a spinning quantum black hole? While our derivation of equation (39), based on the entropy of Hawking radiation was for a static q-metric, the fact that the dimensionless $q$-parameter changes by $5\%$ across 10 orders of magnitude in mass, suggests that it may also be insensitive to BH spin. More conservatively, given that $a_{*}\to 0$ is a continuous limit of Kerr spacetime, we may speculate that $a_{*}\sim{\cal O}(1)$ would only yield ${\cal O}(1)$ changes in $q$, thus keeping it within 102010-20% range (at least away from the extremal limit).
Figure 1: (Top, a) Predicted dependence of minimum $q$ parameter on the mass and spin of astrophysical black holes (equation 39). (Bottom, b) The range of quantum corrections induced in the multipolar structure of the Kerr metric, for the same span of black hole masses/spins.
With this preamble, equation (39) will be used to forecast anticipated departures from Kerr geometry, due to quantum violations of the no-hair theorem, within the context of $(Q_{1},Q_{2})$ parameters. However, it is important to note that the precise dependence on spin should not be taken too seriously. Fig. (1a) shows the values of $|q|$ predicted from equation (39) for astrophysical black hole masses, while Fig. (1b) plots the resulting $Q_{1}$ and $Q_{2}$ parameters. For astrophysical entropies $S\,\sim\,10^{78}\,-\,10^{97}$ the microstate bound implies $q\simeq 0.2$, well within the regime where the series converges and small enough that the metric still approximates Schwarzschild outside a few gravitational radii; we shall come back to this point below.
V.2 Observational signatures
With these predictions at hand, we shall next explore how two forthcoming astrophysical experiments can probe these quantum violations of the no-hair theorem.
V.2.1 Next Generation Event Horizon Telescope (ngEHT)
At (sub)millimetre wavelengths, the brightness distribution around a supermassive black hole encodes strong lensing geometry at radii $r\sim\mathcal{O}(m)$. The most robust geometric observables are: (i) the photon ring diameter and its departures from circularity (e.g., axial ratio, quadrupolar distortions) and (ii) the image centroid (or ring centre) relative to the dynamical centre. These quantities depend primarily on the spacetime lensing map and are comparatively insensitive to the detailed emissivity once multifrequency coverage is used.
[…]
V.2.2 Laser Interferometer Space Antenna (LISA)
Let us now turn our focus to potential observations of extreme-mass-ratio inspirals (EMRIs) by the space-based Laser Interferometer Space Antenna (LISA). In an EMRI, a compact object of mass $\mu\,\ll\,m$ orbits a massive black hole and slowly inspirals through $\sim 10^{4}-10^{5}$ cycles in band e.g., PhysRevD.95.103012 . The phasing and precession structure of the gravitational waveform encode the multipole structure of the central object. In Kerr geometry, one has $Q_{2}=-Ma^{2}$; any deviation of the mass quadrupole $Q_{2}$ from this relation accumulates a measurable phase shift over the radiation-reaction time. Of course, spin and quadrupole both affect precession, so one needs at least one additional independent multipolar handle to break the quadrupole-spin degeneracy, practically, this comes from combining multiple observables: the two precessions and the frequency‑dependent sideband structure (inclination/polar motion harmonics) and late inspiral behaviour near the last stable orbit. In our language, this is equivalent to measuring one extra Kerr-null combination beyond the quadrupole, so that the fit cannot trade spin against $q$. With LISA‑class signal‑to‑noise, long coherent baselines, and waveform models that include precession and higher harmonics, a distortion as large as $q\sim 0.2$ would be potentially identifiable PhysRevD.52.5707 ; PhysRevD.75.042003 ; PhysRevD.95.103012 ; 2023LRR….26….2A .
VI Summary and conclusions
We have shown that a single geometric deformation; the small, positive static quadrupole encoded by the $q$-metric removes the ultraviolet divergence of the entropy of Hawking radiation that plagues the brick wall treatment of Schwarzschild. In the extreme near-horizon region the angular metric acquires anisotropic scaling, $g_{\theta\theta}\propto\rho^{2q}$ and $g_{\phi\phi}\propto\rho^{-2q}$, which modifies the centrifugal barrier and converts the Schwarzschild logarithm into a convergent power. Counting quasinormal modes in the WKB approximation and performing both the radial overtone and angular sums in closed form, we obtained a compact expression for the canonical entropy that tends to the usual logarithmic divergence in the limit $q$ approaching zero.
Combining that microscopic result with the macroscopic requirement that the entropy trapped within the angular momentum barrier does not exceed the Bekenstein-Hawking value sets the lower bound of $|q|\simeq 0.18$-$0.19$ for astrophysical black holes. In other words, the distortion that cures the brick wall is not arbitrarily small: thermodynamic self-consistency fixes its natural scale.
Although our derivation used the static $q$-metric, the same parameter controls clean diagnostics in a stationary extension. In particular, multipole combinations that vanish for Kerr acquire sensitivity to $q$ (thus violating the no-hair theorem), while near-horizon lensing observables (photon ring shape and centroid) are perturbed at the percent-to-ten percent level for $|q|\,\sim\,0.2$. These signatures are natural targets for the ngEHT and gravitational wave spectroscopy of EMRIs with LISA. In short, observationally testable quadrupole distortion thus links ultraviolet finiteness, gravitational wave echo phenomenology, and forthcoming observations of supermassive black holes, offering a concrete bridge between horizon-scale quantum physics and astronomical measurements. On the theory front, extending the analysis to spinning backgrounds and/or additional field content, as well as quantifying backreaction and waveform systematics for LISA, are natural next steps.
Acknowledgements.
We would like to thank Suvendu Giri, Naritaka Oshita, and Michael Florian Wondrak for useful discussions. This research is supported in part by the University of Waterloo, Natural Sciences and Engineering Research Council of Canada, by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities at Perimeter Institute.
Appendix A Klein-Gordon equation components
Time derivative
$\frac{1}{\sqrt{-g}}\partial_{t}\,\bigl{(}\sqrt{-g}\,g^{tt}\partial_{t}\phi\bigr{)}=-\frac{1}{f(\rho)}\,\partial_{t}^{2}\phi.$ (43)
Radial derivative
$\frac{1}{\sqrt{-g}}\,\partial_{\rho}\bigl{(}\sqrt{-g}\,g^{\rho\rho}\partial_{\rho}\phi\bigr{)}$ $=\frac{1}{4m^{2}\sqrt{f}\sin\theta}\,\partial_{\rho}\Bigl{(}4m^{2}\sqrt{f}\sin\theta\,\partial_{\rho}\phi\Bigr{)}$ (44)
$=\partial_{\rho}^{2}\phi+\tfrac{1}{2}\,\frac{f^{\prime}(\rho)}{f(\rho)}\,\partial_{\rho}\phi.$
Angular $\phi$ term
$\frac{1}{\sqrt{-g}}\,\partial_{\phi}\bigl{(}\sqrt{-g}\,g^{\phi\phi}\partial_{\phi}\phi\bigr{)}$ $=\frac{1}{4m^{2}\sqrt{f}\,\sin\theta}\,$ (45)
$\times\partial_{\phi}\!\left(4m^{2}\sqrt{f}\sin\theta\,\frac{1}{4m^{2}}\left(\frac{\rho}{4m}\right)^{2q}\frac{1}{\sin^{2}\theta}\,\partial_{\phi}\phi\right)$
$=\frac{1}{4m^{2}}\,\left(\frac{\rho}{4m}\right)^{2q}\,\frac{1}{\sin^{2}\theta}\,\partial_{\phi}^{2}\phi.$
Angular $\theta$ term
$\frac{1}{\sqrt{-g}}\,\partial_{\theta}\bigl{(}\sqrt{-g}\,g^{\theta\theta}\partial_{\theta}\phi\bigr{)}$ $=\frac{1}{4m^{2}\sqrt{f}\,\sin\theta}\,$ (46)
$\times\partial_{\theta}\!\left(4m^{2}\sqrt{f}\sin\theta\,\frac{1}{4m^{2}}\left(\frac{\rho}{4m}\right)^{-2q}\partial_{\theta}\phi\right)$
$=\frac{1}{4m^{2}}\,\frac{1}{\sin\theta}\,\partial_{\theta}\!\left(\sin\theta\left(\frac{\rho}{4m}\right)^{-2q}\partial_{\theta}\phi\right).$
Adding the four pieces gives equation (10) as
$\Box\varphi$ $\,=-\frac{1}{f(\rho)}\,\partial_{t}^{2}\varphi+\partial_{\rho}^{2}\varphi+\frac{1}{2}\,\frac{f^{\prime}(\rho)}{f(\rho)}\,\partial_{\rho}\varphi$ (47)
$+\frac{1}{4m^{2}}\left[\left(\frac{\rho}{4m}\right)^{-2q}\frac{1}{\sin\theta}\,\partial_{\theta}\!\left(\sin\theta\,\partial_{\theta}\varphi\right)+\left(\frac{\rho}{4m}\right)^{2q}\frac{1}{\sin^{2}\theta}\,\partial_{\phi}^{2}\varphi\right].$